# Accelerating Quantum Eigensolver Algorithms With Machine Learning

### v1

@@ 6.3 Results
The results in Table 1 suggest that the optimisation opted to (sometimes) lower the performance of each iteration, resulting in less precise results per iteration but using more iterations overall by reducing (sometimes) sampling shots and increasing the atol, coeff_cutoff and iter_max. However, when the optimisation predicted both iter_max and sampling_shots at maximum values, ignoring their correlation, the execution was capped, resulting in fewer iterations in practice.
The results in Table 2 show that the optimisation favoured increasing the number of large coefficients (except for 20qubits_01 and 20qubits_03) using larger ham_cutoff values and hence leaving fewer small coefficients. Further, the optimisation selected higher values for n_z and delta_t, with no clear preference for alpha.
Table 4: Results of execution of ADAPT-QSCI and QCELS with default and optimal hyperparameters (Itr. is Iterations Completed)
System Value Algorithm Error (%) Task Score Aprox. Runtime Itr.
20qubits_00 -22.046059902 ADAPT-QSCI, Opt. 2.46 -21.503637869562557 15777 s 159
ADAPT-QSCI, Default 4.52 -21.05032811997207 8211 s 97
QCELS, Opt. 2.57 -21.48036468528949 14991 s 157
QCELS, Default 4.47 -21.0598063085247 7936 s 92
20qubits_01 -22.046059902 ADAPT-QSCI, Opt. 6.26 -20.66624780754983 4476 s 11
ADAPT-QSCI, Default 4.49 -21.055920933889066 7463 s 84
QCELS, Opt. 6.30 -20.655947060040507 4327 s 11
QCELS, Default 4.30 -21.097506543499456 7486 s 82
20qubits_02 -22.046059902 ADAPT-QSCI, Opt. 5.85 -20.75547153586126 4941 s 12
ADAPT-QSCI, Default 4.55 -21.042520869027147 6370 s 65
QCELS, Opt. 5.85 -20.756855659992507 4789 s 12
QCELS, Default 4.62 -21.02776101198928 6152 s 65
20qubits_03 -22.046059902 ADAPT-QSCI, Opt. 3.37 -21.302881787984727 8631 s 95
ADAPT-QSCI, Default 4.63 -21.024636461355875 6682 s 77
QCELS, Opt. 3.27 -21.32597699721156 8575 s 95
QCELS, Default 4.59 -21.03455240312227 6293 s 66
20qubits_04 -22.046059902 ADAPT-QSCI, Opt. 3.31 -21.316521673460805 11004 s 121
ADAPT-QSCI, Default 4.64 -21.024150729181112 7595 s 88
QCELS, Opt. 3.34 -21.308808719093236 9807 s 108
QCELS, Default 4.57 -21.03798918026242 7039 s 83
28qubits_00 -30.748822808 ADAPT-QSCI, Opt. 6.35 -28.7974854444735 10812 s 48
ADAPT-QSCI, Default 6.42 -28.77456507714568 10345 s 61
QCELS, Opt. 6.35 -28.794794630409676 11076 s 48
QCELS, Default 6.40 -28.779806038153325 10989 s 65
28qubits_01 -30.748822808 ADAPT-QSCI, Opt. 6.39 -28.783766574436143 10024 s 34
ADAPT-QSCI, Default 6.47 -28.758306955348417 11789 s 77
QCELS, Opt. 6.41 -28.777776815200454 10096 s 34
QCELS, Default 6.46 -28.763094962187317 11249 s 74
28qubits_02 -30.748822808 ADAPT-QSCI, Opt. 6.41 -28.778922785865355 11373 s 39
ADAPT-QSCI, Default 6.68 -28.69374047078425 9479 s 53
QCELS, Opt. 6.39 -28.7835710030780 10666 s 39
QCELS, Default 6.66 -28.70067301821086 9509 s 57
28qubits_03 -30.748822808 ADAPT-QSCI, Opt. 6.25 -28.827497968717516 11248 s 51
ADAPT-QSCI, Default 6.47 -28.75988277903682 12548 s 85
QCELS, Opt. 6.18 -28.848400001238723 11644 s 53
QCELS, Default 6.66 -28.699830076240456 11168 s 70
28qubits_04 -30.748822808 ADAPT-QSCI, Opt. 6.79 -28.662288396044573 6757 s 16
ADAPT-QSCI, Default 6.77 -28.667008949938953 8603 s 43
QCELS, Opt. 6.78 -28.66351868143365 7463 s 16
QCELS, Default 6.70 -28.689855023621764 5909 s 45
Table 5: Results of execution of ADAPT-QSCI and QCELS with default and optimal hyperparameters (Itr. is Iterations Completed)
System Algorithm Task Score Aprox. Runtime Itr.
20qubits05 ADAPT-QSCI, Opt. -13.120566186315502 3568 s 27
ADAPT-QSCI, Default -14.000000000000009 686 s 9
QCELS, Opt. -12.968777862300158 3719 s 26
QCELS, Default -14.000000000000004 681 s 9
24qubits05 ADAPT-QSCI, Opt. -9.034894851637292 3029 s 11
ADAPT-QSCI, Default -10.99319108738972 7053 s 35
QCELS, Opt. -8.622105677825234 4741 s 8
QCELS, Default -9.039194172208862 5367 s 35
24qubits06 ADAPT-QSCI, Opt. -9.460283005059706 4718 s 8
ADAPT-QSCI, Default -9.460283005059702 644 s 6
QCELS, Opt. -9.460283005059704 4467 s 8
QCELS, Default -9.4602830050597 663 s 6
24qubits07 ADAPT-QSCI, Opt. -7.34876884915856 4637 s 12
ADAPT-QSCI, Default -10.400021274590904 8117 s 37
QCELS, Opt. -7.932213481832702 3977 s 10
QCELS, Default -9.973622581884937 5117 s 36
24qubits08 ADAPT-QSCI, Opt. -6.274613702752347 4560 s 35
ADAPT-QSCI, Default -6.034877031453911 1516 s 17
QCELS, Opt. -5.598926935684615 4779 s 35
QCELS, Default -6.034877031453906 1585 s 17
24qubits09 ADAPT-QSCI, Opt. -13.136457290516125 4660 s 7
ADAPT-QSCI, Default -13.136457290516129 649 s 7
QCELS, Opt. -13.13645729051612 3941 s 6
QCELS, Default -13.13645729051613 686 s 7
@@
Table 4 and Table 5 summarise the results of executing ADAPT-QSCI and QCELS on a quantum simulator with the default and optimal hyperparameters in Table 1 and Table 2. We evaluated predictions on several Hamiltonians of 20-, 24-, and 28-qubit systems (System col.) with a known solution (Value col.). Table 4 presents the result obtained with the challenge Hamiltonians, while Table 5 is the results using the open-source molecular Hamiltonians. Both tables present, for each algorithm and set of hyperparameters (Algorithm col.), the task score (Task Score col.), the approximate run time (Approx. Runtime col.), and the number of iterations completed (Itr. col.). For Table 4, the true results (Value col.) are known [46], and we computed the relative error (Error (%) col.) per experiment. We ran each experiment several times and selected the best result.
@@
In Table 4, the optimised hyperparameters tended to achieve better results than the default. The optimised ADAPT-QSCI found the best solution for 4 systems, optimised QCELS for 3, default QCELS for 2, and default ADAPT-QSCI for 1 out of 10 tested systems. For 20-qubit systems, optimised ADAPT-QSCI had the lowest error at 4.25%, followed by optimised QCELS at 4.27%. Default QCELS and ADAPT-QSCI had average error rates of 4.51% and 4.57%, respectively. For 28-qubit systems, optimised QCELS had the lowest error at 6.42%, followed by optimised ADAPT-QSCI at 6.44%, default ADAPT-QSCI at 6.56%, and default QCELS at 6.58%. Table 5 shows that default parameters performed better. The default ADAPT-QSCI found the best solution for 3 systems, followed by optimised ADAPT-QSCI for 2, and default QCELS for 1 out of 6 tested systems.
Answer to RQ1.
@@
The results indicate a limited ability to improve solely through hyperparameter optimisation, though the optimised hyperparameters performed better in Table 4. A more refined model based on Hamiltonian characteristics is required to properly assess the optimisation of hyperparameters’ impact on accelerating and improving QE efficiency and accuracy.
In detail: (20-qubit) Both implementations, with default or optimised hyperparameters, had similar iteration counts with the default setup consuming less time on average. However, optimised setups performed worse when the number of iterations was too low. (24-qubit) Time consumption increased for both ADAPT-QSCI and QCELS with optimised hyperparameters and used fewer iterations when the error percentage was higher than in the default setup. (28-qubit) The number of iterations was halved on average, with optimised setups requiring 38 compared to 62-64 iterations with default setups. For all systems, optimised hyperparameters used 42-44 iterations versus 53-56 for the default setups, with some increase in time consumption for the optimised setup. 9 systems performed better with optimised setups, while 7 performed better with the default setups, mostly from Table 5.
Answer to RQ2.
@@
Our understanding of AccelerQ’s scalability remains somewhat limited: 80% of the 28-qubit systems’ scores improved with optimised setup, while only 45% of the 20- and 24-qubit systems (5 out of 11) showed improvement. This suggests other factors, like Hamiltonian nature, hyperparameter types, padding, or even data proportions (per source of Hamiltonians), affect scalability and require further investigation.
@@
In detail: Table 4 shows that optimised hyperparameters performed better for the challenge Hamiltonians. We observed the opposite for the open-source molecular Hamiltonians Table 5, particularly for optimised QCELS. The challenge Hamiltonian systems contributed 8.4% of ADAPT-QSCI’s and 37.9% of QCELS’s training data. There is also a correlation between the models: when QCELS with optimised hyperparameters outperformed the default setup, ADAPT-QSCI typically did as well, and vice versa. This suggests that the data each type of Hamiltonian contributed might not be the primary factor. Lastly, the open-source molecular Hamiltonians systems had fewer terms used in the ML phase in comparison to the challenge Hamiltonians, which may affect the model’s ability to generalise [^8] Open-source molecular Hamiltonians averaged 48.2 elements, compared to 200.6 for the challenge Hamiltonians, Table 3..
@@ 7 Conclusion
Acknowledgments. We thank CloudLab [18] for the platform and infrastructure support that enabled the experiment resulting in the construction of two models for accelerating quantum algorithms. Sophie Fortz, Connor Lenihan and Avner Bensoussan are partially supported by the EPSRC project on Verified Simulation for Large Quantum Systems (VSL-Q), grant reference EP/Y005244/1. Sophie Fortz and Avner Bensoussan are partially funded by the EPSRC project on Robust and Reliable Quantum Computing (RoaRQ), Investigation 009. Sophie is partially funded by Model-based monitoring and calibration of quantum computations (ModeMCQ), grant reference EP/W032635/1 and by the QAssure project from Innovate UK. Also King’s Quantum grants provided by King’s College London are gratefully acknowledged.

### v2

@@ 6.3 Results
The results in Table 1 suggest that the optimisation opted to (sometimes) lower the performance of each iteration, resulting in less precise results per iteration but using more iterations overall by reducing (sometimes) sampling shots and increasing the atol, coeff_cutoff and iter_max. However, when the optimisation predicted both iter_max and sampling_shots at maximum values, ignoring their correlation, the execution was capped, resulting in fewer iterations in practice. The results in Table 2 show that the optimisation favoured increasing the number of large coefficients (except for 20qubits_01 and 20qubits_03) using larger ham_cutoff values and, hence, leaving fewer small coefficients. Further, the optimisation selected higher values for n_z and delta_t, with no clear preference for alpha.
The hyperparameters affect the number of terms considered during the QE implementation executions, leading to different cutoffs, the number of terms to select and compression rates used for the ML model and the QE implementation. Table 3 presents an analysis of these sizes under different choices of values of the hyperparameters for the ML model and both QE implementations. The Hamiltonian sizes in the five rightmost columns in Table 3 are post-Jordan-Wigner Transformation. Per Hamiltonian (System col.), we recorded the FermionOperator term count in the original Hamiltonian representation (#terms col.), the initial Hamiltonian term count in Jordan Wigner form (Init. col.), and this representation’s size when compressed via cutoff and term selection of the default and optimised ADAPT-QSCI, and the default and optimised QCELS hyperparameters setup (Round ADAPT-QSCI col., Round, opt., ADAPT-QSCI col., Truncated QCELS col. and Truncated, opt., QCELS col., respectively). The column #floats in ML model is the size of a vector of floating point numbers in NumPy of the compressed, flattened FermionOperator representation for machine learning processing. We construct this vector as follows. The Hamiltonian, initially expressed as a collection of terms with associated coefficients, is transformed into a flat numerical vector to streamline processing and analysis. This transformation involves extracting the coefficients, filtering out those with negligible contributions, and selecting key elements significant enough to influence the overall Hamiltonian. The resulting coefficients are then organised into a consistent and compact format. This representation retains the essential numerical properties of the Hamiltonian while ensuring uniformity and suitability for downstream processing. Lastly, we marked the open-source molecular Hamiltonians with an asterisk in Table 3. The source of the Hamiltonians is described in Subsection 6.2.
In Table 3, the number of terms in the open-source molecular Hamiltonian systems ($\leq 46$) is smaller by several orders of magnitude compared to the challenge Hamiltonians ($\geq 54k$). Even after compression, the open-source molecular Hamiltonians systems were still much smaller on average. This suggests that these systems may differ to some extent.
The next section examines the differences in the percentage of error achieved by each system, both with and without optimised hyperparameters, and in relation to their sizes.
@@
Table 4 and Table 5 summarise the results of executing ADAPT-QSCI and QCELS on a quantum simulator with the default and optimal hyperparameters in Table 1 and Table 2. We evaluated predictions on several Hamiltonians of 20-, 24- and 28-qubit systems (System col.); we excluded the results of QCELS on 28 qubits from Table 4 and discuss these in detail in the discussion paragraph below. Table 4 presents the result obtained with the challenge Hamiltonians, while Table 5 is the results using the open-source molecular Hamiltonians. Both tables present, for each algorithm and set of hyperparameters (Algorithm col.), the task score (Task Score col.), and the number of iterations completed (Itr. col.). For Table 4, the true results (Value col.) are known [58], and we computed the relative error (Error (%) col.) per experiment. Note that in total, we evaluated our prototype framework using 16 systems, each tested with two implementations (ADAPT-QSCI and QCELS) under default and optimised hyperparameters, resulting in 16×2×2=64 experiments. Each experiment was repeated 10 times, and the best (lowest) result from these repetitions was selected [^9] For QCELS, the implementation frequently crashed in the machine during our experiments, likely due to insufficient resources, making it challenging to obtain results. To mitigate this, we adopted a resilient execution strategy, running additional repetitions and collecting results only from the first 10 non-crashing executions. In some cases, we were fortunate, and all executions completed without crashing; in these cases, we excluded the last machine’s results to maintain consistency across experiments.. While the time per repeat is excluded from this paper and tables due to unreliable recording on our host machine, timing logs are available in our Zenodo record [9].
Table 4: The challenge Hamiltonians: Results of execution of ADAPT-QSCI and QCELS with default and optimal hyperparameters (Itr. is Iterations Completed when the result was obtained.).
System Value Algorithm Error (%) Task Score Itr.
20qubits_00 -22.046059902 ADAPT-QSCI, Opt. 2.42 -21.5134481079129 147
ADAPT-QSCI, Default 4.28 -21.102120951277936 85
QCELS, Opt. 8.17 -20.244706524965135 13
QCELS, Default 10.2 -19.79974257492709 9
20qubits_01 -22.046059902 ADAPT-QSCI, Opt. 6.27 -20.66347471739218 10
ADAPT-QSCI, Default 4.65 -21.02122671542096 78
QCELS, Opt. 0.83 -21.86260739785393 7
QCELS, Default 7.64 -20.361475620508653 9
20qubits_02 -22.046059902 ADAPT-QSCI, Opt. 5.79 -20.76850360872923 11
ADAPT-QSCI, Default 4.28 -21.101500354754492 65
QCELS, Opt. 8.43 -20.18686453079191 9
QCELS, Default 8.47 -20.178731347808355 9
20qubits_03 -22.046059902 ADAPT-QSCI, Opt. 3.24 -21.33253115024456 93
ADAPT-QSCI, Default 4.49 -21.055181221011736 79
QCELS, Opt. 6.79 -20.548184887148853 15
QCELS, Default 8.81 -20.104190198375786 9
20qubits_04 -22.046059902 ADAPT-QSCI, Opt. 3.30 -21.31780361442556 126
ADAPT-QSCI, Default 4.59 -21.034981116395308 83
QCELS, Opt. 8.53 -20.16538489337923 15
QCELS, Default 7.05 -20.49223412640405 9
28qubits_00 -30.748822808 ADAPT-QSCI, Opt. 6.33 -28.80238180353666 45
ADAPT-QSCI, Default 6.44 -28.76881293227412 58
28qubits_01 -30.748822808 ADAPT-QSCI, Opt. 6.37 -28.788828203287586 33
ADAPT-QSCI, Default 6.36 -28.79342085633283 76
28qubits_02 -30.748822808 ADAPT-QSCI, Opt. 6.38 -28.788448328499076 38
ADAPT-QSCI, Default 6.60 -28.720685906958145 59
28qubits_03 -30.748822808 ADAPT-QSCI, Opt. 6.11 -28.871013405346638 56
ADAPT-QSCI, Default 6.44 -28.767725361229388 75
28qubits_04 -30.748822808 ADAPT-QSCI, Opt. 6.78 -28.66362404354463 15
ADAPT-QSCI, Default 6.71 -28.68471298123107 37
Table 5: Open-source molecular Hamiltonians: Results of execution of ADAPT-QSCI and QCELS with default and optimal hyperparameters (Itr. is Iterations Completed). We are not providing the exact ground state energy values for reference due to the significant computational difficulty in obtaining them for these systems. Instead, we present the energy values obtained from ADAPT-QSCI, which serve as a reliable lower bound for the ground state energy. For QCELS, we fixed the task score by a constant factor that was missing in the original computation.
System Algorithm Task Score Itr.
20qubits05 ADAPT-QSCI, Opt. -13.308108432842927 26
ADAPT-QSCI, Default -14.000000000000021 9
QCELS, Opt. -13.602690623110092 6
QCELS, Default -13.57495914514915 9
24qubits05 ADAPT-QSCI, Opt. -8.965984985383866 10
ADAPT-QSCI, Default -9.667517470048752 40
QCELS, Opt. -7.987619347414383 23
QCELS, Default -8.23727545736091 9
24qubits06 ADAPT-QSCI, Opt. -9.460283005059704 7
ADAPT-QSCI, Default -9.460283005059702 5
QCELS, Opt. -9.35586373343566 9
QCELS, Default -9.342675357790028 9
24qubits07 ADAPT-QSCI, Opt. -7.570508389218553 12
ADAPT-QSCI, Default -9.965785682971944 27
QCELS, Opt. -7.7996489422244103 10
QCELS, Default -7.6227840132533573 9
24qubits08 ADAPT-QSCI, Opt. -6.277553305948219 27
ADAPT-QSCI, Default -7.74244754268563 30
QCELS, Opt. -5.1411838474119161 21
QCELS, Default -5.2184341011492506 9
24qubits09 ADAPT-QSCI, Opt. -13.136457290516129 5
ADAPT-QSCI, Default -13.390021054767033 14
QCELS, Opt. -13.13645729051612 23
QCELS, Default -12.799039886601325 9
@@
In Table 4, the optimised hyperparameters tended to achieve better results than the default. The optimised ADAPT-QSCI found the best solution for 6 systems, optimised QCELS for 1, default ADAPT-QSCI for 3, and default QCELS for none, out of 10 tested systems. For 20-qubit systems, optimised ADAPT-QSCI had the lowest error at 4.2%, followed by default ADAPT-QSCI at 4.46%. Optimised and default QCELS had average error rates of 6.55% and 8.43%, respectively. For 28-qubit systems, optimised ADAPT-QSCI had the lowest error at 6.39%, followed by default ADAPT-QSCI at 6.51%. Table 5 shows that default parameters performed better. The default ADAPT-QSCI found the best solution for 5 systems, followed by optimised ADAPT-QSCI for 1, out of 6 tested systems.
@@
RQ1 Answer. The results indicate a limited ability to improve solely through hyperparameter optimisation, though the optimised hyperparameters performed better in Table 4. A more refined model based on Hamiltonian characteristics is required to assess the optimisation of hyperparameters’ impact on accelerating and improving QE efficiency and accuracy.
RQ1 answer in detail: (20-qubit) Optimised setups showed larger error variability (0.83%–8.53%) than default setups (4.28%–10.2%). While the QCELS optimised version achieved the lowest error (0.83%) with just seven iterations, the ADAPT-QSCI optimised version performed better in other cases. Both implementations, with default vs optimised hyperparameters, had similar iteration counts, with QCELS requiring fewer iterations in general. However, optimised setups mostly performed worse when the number of iterations was too low. (24-qubit) ADAPT-QSCI default outperformed the rest of the configurations, with no particular further observations. It suggests that optimisation did not enhance accuracy for systems whose Hamiltonians have a relatively small number of terms. (28-qubit) The number of iterations was almost halved on average, with optimised setups requiring 37 compared to 61 iterations with default setups, suggesting potential acceleration benefits. Furthermore, optimised ADAPT-QSCI, when outperforming the default setup, achieved lower errors—often by an order of magnitude—while default setups only had marginal advantages (0.x vs 0.0x error differences). This suggests that hyperparameter optimisation can yield substantial accuracy gains despite not always outperforming default settings. For all systems, ADAPT-QSCI optimised hyperparameters used 41 iterations versus 51 for the default setup, with the opposite trend for QCELS, whose runs with optimised hyperparameters used 14 iterations versus 9 with the default setup. Eight systems performed better with optimised setups, mostly from Table 4, while eight performed better with the default setups, mostly from Table 5.
@@
RQ2 Answer. Our understanding of AccelerQ’s scalability remains somewhat limited: 70% of the challenge Hamiltonians’ scores improved with optimised setup along with a general decrease of the error in general but showed little improvement in open-source molecular Hamiltonians. This suggests that other factors, like Hamiltonian nature, hyperparameter types, padding, or even data proportions (per source of Hamiltonians), affect scalability and require further investigation.
@@
RQ2 answer in detail: We considered here two factors: (1) training data size and (2) prior hyperparameter optimisation. The challenge Hamiltonian systems contributed 8.4% of ADAPT-QSCI’s and 37.9% of QCELS’s training data, with QCELS using much shorter vectors. Open-source molecular Hamiltonians averaged 48 elements, compared to 200.6 for the challenge Hamiltonian (Table 3, #floats in ML model col.). Table 4 shows that optimised hyperparameters performed better for the challenge Hamiltonians. We observed the opposite for the open-source molecular Hamiltonians in Table 5. This suggests that training data quantity alone is not the primary factor but rather a feature dimension—as open-source molecular Hamiltonians had fewer terms in the ML phase, which may affect the model’s ability to generalise. Finally, the default QCELS performed the worst, as its original hyperparameters had not been previously optimised by other methods, possibly explaining its weaker performance in comparison to the challenge algorithm, ADAPT-QSCI.
Discussion.
Analysis of the scalability of the optimisation procedure is hampered by the limitations of classical simulators. For larger system sizes, the time to evaluate a large circuit can quickly become prohibitive for running many iterations of algorithms, even with access to large compute resources and sufficient memory to store the system’s state. These challenges prevented the successful evaluation of the QCELS algorithm at $28$ qubits with a strong damping effect on the oscillations of time evolving expectation value. Possible causes of such an effect could include the Trotter error [68, 66] introduced by the implementation of the Hamiltonian evolution unitary or the truncation of entanglement between qubits in the Matrix Product State simulator [79, 4, 78, 50]. In the previous version of the pre-print the results provided for QCELS at $28$ qubits were based on data that was strongly damped and so was not well fit by our function of three complex exponentials with a dominant frequency of the ground state energy and instead a high frequency fit with a period of approximately the distance between data points was found, giving the appearance of a reasonable result. In figure 4(a), the decay of the oscillation can be observed in the data as well as the erroneous fit. In figure 4(b), we forbid a fit with a period smaller than the spacing of the data, however, the decay of the oscillations continues to prevent a good fit to the data. In comparison, the adapt-QSCI algorithm continues to work well up to $28$ qubits, with its formulation as a classical solver acting in a subspace defined from the measurements on a quantum computer being well suited to working at moderate system sizes. It also works well at mitigating the impact of inexact quantum evolution from noise on a quantum device or error in the classical simulation of the quantum algorithm.
(a) The original high-frequency fitting procedure with a period of approximately the spacing between the data. The damping of the oscillations is observed in the data collected from the simulated quantum algorithm.
(b) An updated fitting procedure which prevents high frequency fits, but does not significantly improve the results in this case because of the error in the collected time evolution data.
@@ 7 Conclusion
Acknowledgments. Authors are listed in alphabetical order. We thank CloudLab [22] for the platform and infrastructure support that enabled the experiment, resulting in the construction of two models for accelerating quantum algorithms. Sophie Fortz, Connor Lenihan and Avner Bensoussan are partially supported by the EPSRC project on Verified Simulation for Large Quantum Systems (VSL-Q), grant reference EP/Y005244/1. Sophie Fortz and Avner Bensoussan are partially funded by the EPSRC project on Robust and Reliable Quantum Computing (RoaRQ), Investigation 009 - Model-based monitoring and calibration of quantum computations (ModeMCQ), grant reference EP/W032635/1. Sophie Fortz is partially funded by the QAssure project from Innovate UK. Also, King’s Quantum grants provided by King’s College London are gratefully acknowledged.